# Reconfigurable generation of spatial entanglement in nonlinear waveguide arrays

### v1

@@ Abstract
Yet, an additional layer of complexity can introduce even larger possibilities, by adding to the continuous nature of the interference process the ability to continuously generate photons all along the device. This can be realized in a waveguide array made out of a nonlinear material, by injecting a classical pump beam generating stochastically photon pairs by spontaneous parametric down-conversion (SPDC) or four-wave mixing [27, 28, 29, 30]. This approach offers novel configurations that have no equivalent in bulk optics nor discrete photonic circuits due to the intricate combination of the generation and manipulation steps of photonic quantum states [31, 32, 33].
@@
Figure 1: (a) Principle of the spatial entanglement engineering in an AlGaAs nonlinear waveguide array. Here, two continuous-wave pump laser beams (sketched in red) with a phase offset $\varphi$ between them generate photon pairs that undergo quantum walks (blue), resulting in spatial entanglement over the whole lattice. (b) SEM image of a fabricated sample, showing a cut of the waveguide array in the central part. (c) Wider-field SEM image showing the array connected to S-bend waveguides for the individual optical addressing of each spatial channel.
@@
In this work, we exploit this concept in AlGaAs-based nonlinear waveguide arrays to implement a compact and versatile source of spatially entangled states of light, operating at room temperature and telecom wavelength. We use a double-pump configuration to engineer the output quantum state and implement different types of spatial correlations, violating a non-classicality criterion by several tens of standard deviations. This control relies on a quantum interference effect between the biphoton wavefunction generated in each pumped waveguide, which is fully accounted for by analytical calculations and numerical simulations. Coupled with the high integration capabilities of the AlGaAs platform [34, 35] and the ability to implement diverse lattice geometries, this demonstration paves the way to quantum information tasks leveraging the high-dimensional spatial degree of freedom in continuously-coupled photonic systems.
@@
The working principle of our AlGaAs nonlinear waveguide array is sketched in Fig. 1a. A classical pump beam (wavelength $\simeq 775$ nm, sketched in red) injected into one or several waveguides generates photon pairs at telecom wavelength ($\simeq 1550$ nm, shown in blue) by SPDC, thanks to the strong $\chi^{(2)}$ nonlinearity of the material. These photon pairs can continuously tunnel from one waveguide to the other during their propagation, implementing random quantum walks. Compared to quantum walks in passive circuits [18, 19, 20], the walkers are here generated directly inside the device and the generation can take place at any position along the propagation axis. Besides a gain of integration, this configuration allows for a significantly higher level of spatial entanglement, due to the interference between quantum walks initiated at all possible longitudinal positions [27]. The device is designed so that the transverse coupling of the pump beam is negligible (which is naturally favored by its twice shorter wavelength), so that the pump remains confined in the initially pumped waveguide(s) during propagation.
The evanescent coupling of the telecom (SPDC) modes in the array introduces a discretized dispersion relation (band structure), modifying the propagation constant in the longitudinal direction. For a given mode polarization, it reads
$\beta(\omega,k)=\beta^{(0)}(\omega)+2\,C(\omega)\,\text{cos}(k)$ (1)
where $\beta^{(0)}(\omega)=\omega\,n_{\rm{eff}}(\omega)/c$ is the propagation constant of a single waveguide at frequency $\omega$, governed by the effective refractive index $n_{\rm{eff}}(\omega)$. The second term, where $C$ is the evanescent coupling constant and $k$ is the dimensionless transverse wavevector (normalized to $1/a$, where $a$ is the lattice periodicity), which defines the phase difference between the neighboring waveguides, corresponds to the modification of the dispersion relation induced by the presence of the array. For a given $\omega$, Eq. (1) is analogous to the dispersion relation of an electron in a periodic potential, as obtained in the tight-binding approximation, and the $\beta(\omega,k)$ correspond to the Bloch modes of the lattice.
After SPDC generation in the array under a monochromatic pump beam (frequency $\omega_{p}$), and filtering the signal and idler around degeneracy (i.e. with frequencies $\omega_{s}\simeq\omega_{i}\simeq\omega_{p}/2$), the biphoton state at the output can be written as a superposition of spatial modes, governed by the complex-valued function
$\phi_{k}(k_{s},k_{i})=\widetilde{A}_{p}(k_{s}+k_{i})\,\operatorname{sinc}\left(\Delta\beta(k_{s},k_{i})\,L/2\right)e^{i\Delta\beta(k_{s},k_{i})\frac{L}{2}}$ (2)
which determines the probability to generate, from an input pump photon, a photon pair with the signal and idler in Bloch modes $k_{s}$ and $k_{i}$ respectively [36]. In this expression, $\widetilde{A}_{p}$ is the spatial Fourier transform of the pump field (determined by its amplitude in each waveguide, and written here by taking into account the conservation of the transverse wavevector), and $L$ is the longitudinal length of the array. In order to obtain a significant pair generation rate, the phase-matching condition $\Delta\beta=0$ must be nearly perfectly satisfied, with $\Delta\beta$ given by
$\Delta\beta(k_{s},k_{i})=\Delta\beta^{(0)}-2\,C_{s}\,\text{cos}(k_{s})-2\,C_{i}\,\text{cos}(k_{i})$ (3)
Here, $\Delta\beta^{(0)}$ is the phase-mismatch of a single (uncoupled) waveguide, which depends on the pump frequency, and $C_{s}=C_{s}(\omega_{p}/2)$ and $C_{i}=C_{i}(\omega_{p}/2)$ are the coupling constants of the signal and idler photons in their corresponding polarization. For a given pump frequency, the spatial correlations can thus be controlled by tailoring the spatial profile of the pump beam (term $\widetilde{A}_{p}$ in Eq. (2)), providing the principle for a compact and reconfigurable source of spatially entangled states of light. The spatial correlations in real space are described by a (symmetric) matrix $\Gamma_{n_{s},n_{i}}=|\phi_{n}(n_{s},n_{i})|^{2}$, where $\phi_{n}(n_{s},n_{i})$ is the Fourier transform of $\phi_{k}$. Note that $\Gamma_{n_{s},n_{i}}$ is not normalized to unity, it takes into account the SPDC efficiency and hence is directly proportional to coincidence counts as will be measured experimentally in the following.
@@
The investigated sample consists in an array of 31 coupled AlGaAs waveguides of length $L=2$ mm (Fig. 1b). The waveguides are 1.7 $\mu$m-wide and separated by 450 nm gaps. The 7 central waveguides of the array are connected to fan-in injection waveguides and fan-out collection waveguides to facilitate their individual optical addressing (Fig. 1c). These injection and collection waveguides have a larger width (6 $\mu$m) to shift their nonlinear resonance wavelength (by 10 nm), so that the injected pump laser generates photon pairs only in the central array region. The total longitudinal length of the sample is $4.3$ mm.
The epitaxial structure is made of a stacking of AlGaAs layers with alternating Aluminum concentrations. On each side of the waveguides core, two Bragg mirrors provide both a photonic band-gap confinement for the pump beam in the NIR range (at $\simeq 775$ nm) and total internal reflection claddings for the produced SPDC photons in the telecom range ($\simeq 1550$ nm). Therefore, the pump and SPDC modes are characterized by different dispersion curves, allowing the single-waveguide phase-matching condition $\Delta\beta^{(0)}=0$ to be satisfied in the spectral range of interest [37, 38, 39, 35]. The epitaxial structure is grown by molecular beam epitaxy, and the sample is processed by electron-beam lithography (using a high-resolution HSQ resist) followed by ICP dry etching. The determined propagation losses are of 0.6 cm-1 at telecom wavelength.
@@
The experiments are carried out by injecting a TE-polarized continuous-wave pump laser with a microscope objective into one or several waveguides (depending on the targeted quantum state), to generate by orthogonally polarized signal and idler photons by type-2 SPDC in the array region. These SPDC photons undergo random quantum walks and are collected at the output of the fan-out waveguides using a commercial lensed fiber array connected to superconducting nanowire single-photon detectors (SNSPD). Correlation events are measured by a time tagger: coincidences between distinct waveguides (i.e. off-diagonal points in the spatial correlation matrix $\Gamma_{n_{s},n_{i}}$) are determined directly using this setup, while coincidences within each waveguide (diagonal correlations) are determined using an additional beamsplitter after each waveguide to split the SPDC photons into its two outputs (with 50% probability) before detection.
@@
Figure 2: (a) Biphoton correlations measured in the central waveguide ($\Gamma_{0,0}$, black), first diagonal ($\Gamma_{1,1}+\Gamma_{-1,-1}$, red) and first antidiagonal ($\Gamma_{1,-1}+\Gamma_{-1,1}$, blue) points of the correlation matrix when pumping the $n=0$ waveguide (raw coincidence counts). (b-c-d) Measured correlation matrices when pumping (b) the central waveguide $n=0$, (c) guides $n=0$ and $n=1$ in phase and (d) guides $n=0$ and $n=1$ out of phase. (e) Cross-correlation term $\Gamma_{0,1}$ measured for varying phase offset $\varphi$ between the two pumped waveguides ($n=0$ and $1$); circles show raw coincidence counts with Poissonian error bar, crosses show noise counts, and the black line is a sinusoidal fit. (f-g-h) Numerically simulated correlation matrices corresponding to the experimental cases (b-c-d); the efficiency unit is the same for all three panels and can thus be directly compared to the experiment.
@@
The pump laser is first injected into the central waveguide ($n=0$) of the array, and coincidences are monitored as a function of the laser wavelength for a coupled pump power of $\simeq$ 1 mW. Here and in the subsequent measurements, a bandpass filter is used before photon detection to reduce the emission bandwidth to 16 nm and effectively suppress (up to 1%) any possible coupling between the spatial and spectral properties of the biphoton states [36]. As shown in Fig. 2a, the coincidences in the central waveguide ($\Gamma_{0,0}$, black points) show a resonance centered at $\lambda_{p,0}=783.45$ nm, which corresponds (up to 0.05 nm) to the resonance wavelength which we measured for a single (uncoupled) waveguide, i.e. the pump wavelength that satisfies $\Delta\beta^{(0)}=0$. The diagonal coincidences ($\Gamma_{1,1}+\Gamma_{-1,-1}$, red points) and antidiagonal coincidences ($\Gamma_{1,-1}+\Gamma_{-1,1}$, blue points) display a similar resonant behavior.
We next set the pump wavelength to $\lambda_{p,0}$ and proceed to the measurement of the full correlation matrix within the Hilbert space formed by the 5 central waveguides. The result is shown in Fig. 2b and compared to numerical simulations in Fig. 2f. As anticipated by the results of Fig. 2a, we observe strong and selective correlations along the diagonal and the antidiagonal axes. That is, photons have an enhanced probability to exit the device either through the same waveguide ($n_{s}=n_{i}$, spatial bunching) or through opposite waveguides ($n_{s}=-n_{i}$, spatial antibunching). Indeed, since $\Delta\beta^{(0)}=0$ the phase-matching condition $\Delta\beta\simeq 0$ selects the Bloch modes satisfying $\text{cos}\,(k_{s})+\text{cos}\,(k_{i})\simeq 0$, corresponding to $k_{s}\pm k_{i}\simeq\pm\pi$ [36]. This corresponds to equal or opposite transverse propagation velocities for the twin photons, $v(k_{s})=\pm v(k_{i})$, giving rise to both a bunching and antibunching pattern [27].
The non-classicality of the measured correlations can be quantified by using the criteria derived by Bromberg et al. [16, 18]. For classical light, the diagonal and off-diagonal correlations must satisfy the inequality $\Gamma_{n_{i},n_{s}}\geq\frac{1}{3}\sqrt{\Gamma_{n_{i},n_{i}}\Gamma_{n_{s},n_{s}}}$. For the measured state of Fig. 2b, this inequality is violated (up to 10 standard deviations) by several points of the measured correlation matrix [36], indicating spatial correlations that cannot be generated by classical means. From the total detected coincidences counts and the measured collection and detection efficiency, the generation rate of these quantum states at the chip output is $1.2\times 10^{5}$ Hz/mW of internal pump power, within a nonlinear generation length of 2 mm. The overlap between the experimental ($\Gamma_{n_{s},n_{i}}^{\rm exp}$) and theoretical ($\Gamma_{n_{s},n_{i}}^{\rm th}$) correlations matrices can be quantified using the similarity parameter [18, 19] $S=\big{(}\sum\sqrt{\Gamma_{n_{s},n_{i}}^{\rm exp}\Gamma_{n_{s},n_{i}}^{\rm th}}\big{)}^{2}/\left(\sum\Gamma_{n_{s},n_{i}}^{\rm exp}\sum\Gamma_{n_{s},n_{i}}^{\rm th}\right)$. We obtain $S=98.4\pm 1.1\%$ (without adjustable parameter), illustrating the high fidelity of the source.
@@
We then set $\varphi=\pi$, corresponding to the $n=0$ and $n=1$ waveguides pumped in phase opposition. The measured correlation matrix is shown in Fig. 2d along with simulation results in 2h (similarity $96.9\pm 1.0\%$). In this case, we observe that spatial bunching is selectively enhanced, while the spreading of correlations is comparable to the single-pump configuration. However, the total SPDC signal (integrated over all measured points of the correlation matrix) is stronger (by a factor $\simeq$ 2.8) for $\varphi=\pi$ than for $\varphi=0$, despite the fact that the input pump power is the same. As the transverse coupling of the pump beam is negligible, the two pump beams do not interfere classically and this interference effect is of quantum nature: the biphoton states generated in the two pumped waveguides interfere constructively when $\varphi=\pi$ and destructively when $\varphi=0$, and it can be shown analytically [36] that the total SPDC intensity is modulated as
@@
This intriguing behavior, with a constructive interference occurring for out-of-phase pumping, can be understood intuitively in the following manner. In the single-pump configuration, correlations are mainly located on the diagonal and antidiagonal (Figs. 2b,f). According to coupled-mode theory, a single photon acquires a phase of $\pi/2$ when it tunnels to a neighboring waveguide; hence, starting from its generation in the central waveguide $n=0$, the biphoton state acquires a phase $\pi$ when it moves one step along the diagonal or the antidiagonal, as this requires two single-photon jumps (either on the same or on opposite sides). The phase of the wavefunction thus changes by $\pi$ between each two successive points along the diagonal and antidiagonal. Now, when pumping two neighboring waveguides $n=0$ and $1$, an interference occurs between the biphoton states |$\ket{\Psi}_{n=0}$ and |$\ket{\Psi}_{n=1}$ generated in these two waveguides. When the two pumps are in phase (resp. out of phase), this interference is destructive (resp. constructive) on the diagonal points (which are aligned for the two states) because of the $\pi$ phase alternation structure described above, in good agreement with the experiments and simulations. This interference is only partial since the two biphoton states have different amplitudes on the interfering points; however it plays a determinant role in the total SPDC efficiency since an important part of the signal is concentrated in each pumped waveguide, providing an intuitive argument for the total SPDC intensity being maximum when $\varphi=\pi$, in agreement with Eq. (5). The situation is reversed in the case of antidiagonal points. However, on the antidiagonal passing through $(0,1)$ the two biphoton states interfere with the exact same amplitude, leading to a perfect interference and thus a high spatial spreading of the correlations when $\varphi=0$, in good agreement with Figs. 2c and 2g. This demonstrated quantum interference effect provides the building block for engineering spatially entangled states in continuously-coupled nonlinear waveguides.
@@
In summary, we have demonstrated a compact and versatile source of spatial entanglement based on quantum walks in semiconductor AlGaAs nonlinear waveguide arrays. Tailoring the spatial profile of the pump field allows to reconfigure the output quantum state and implement various types of spatial correlations, well accounted for by analytical calculations and numerical simulations. Building upon these results, a wider zoology of quantum states could be produced by pumping more waveguides, with controlled intensity and phase relation between them. To achieve this in an efficient and scalable manner, on-chip beamsplitters and phase shifters based on the strong electro-optic effect of AlGaAs [40, 35] could be integrated prior the generation stage. Pushing further the integration, the possibility to integrate the pump laser directly within the nonlinear medium [39], to obtain a fully standalone source, constitutes a distinctive asset of the AlGaAs platform compared to dielectric and silicon-based photonic circuits where spatially entangled states have been studied previously [9].
@@
In addition, the strong second-order nonlinearity of AlGaAs could be leveraged to push nonlinear waveguide arrays into the squeezing regime [41, 42]. In line with recent theoretical studies, this approach is expected to enable the generation of highly multimode squeezed states in a flexible manner [31, 32, 33], establishing promising connections with other continuous-variable multimode platforms such as optical frequency combs [43]. Finally, other lattice geometries can be envisaged to investigate various types of physical phenomena: e.g. modulating the coupling distances according to a Fibonacci [44] or Harper [24] sequence would allow investigating topological protection in the quantum regime, while introducing disorder would enable studying the Anderson localization of multi-particle states. Combined with the possibility of engineering losses to investigate non-Hermitian effects [45] or dynamically modulating the array to harness synthetic dimensions [46, 47], these perspectives make the demonstrated platform appealing to simulate in a controlled environment physical problems otherwise difficult to access in condensed-matter systems.

### v2

# Tunable generation of spatial entanglement in nonlinear waveguide arrays

@@ Abstract
Yet, an additional layer of complexity can introduce even larger possibilities, by adding to the continuous nature of the interference process the ability to continuously generate photons all along the device. This can be realized in a waveguide array made out of a nonlinear material, by injecting a classical pump beam generating stochastically photon pairs by spontaneous parametric down-conversion (SPDC) or four-wave mixing [27, 28, 29, 30, 31]. This approach offers novel configurations that have no equivalent in bulk optics nor discrete photonic circuits due to the intricate combination of the generation and manipulation steps of photonic quantum states [32, 33, 34].
@@
Figure 1: (a) Principle of the spatial entanglement engineering in an AlGaAs nonlinear waveguide array. Here, two continuous-wave pump laser beams (sketched in red) with a phase offset $\varphi$ between them generate photon pairs that undergo quantum walks (blue), resulting in spatial entanglement over the whole lattice. (b) SEM image of a fabricated sample, showing a cut of the waveguide array in the central part. (c) Wider-field SEM image of the sample. (d) Simulated correlation matrix for a linear waveguide array injected with a path-entangled state (|$(\ket{00}+\ket{11})/\sqrt{2}$, and (e) for a nonlinear array when pumping waveguides $n=0$ and $1$ in-phase. (f) Corresponding calculated Schmidt number as a function of the propagation length.
@@
In this work, we exploit this concept in AlGaAs-based nonlinear waveguide arrays to implement a compact and versatile source of spatially entangled states of light, operating at room temperature and telecom wavelength. We use a double-pump configuration to engineer the output quantum state and implement different types of spatial correlations, violating a non-classicality criterion by several tens of standard deviations. This control relies on a quantum interference effect between the biphoton wavefunction generated in each pumped waveguide, fully accounted for by analytical calculations and numerical simulations, and paving the way to quantum information tasks leveraging the high-dimensional spatial degree of freedom in continuously-coupled photonic systems.
@@
The working principle of our AlGaAs nonlinear waveguide array is sketched in Fig. 1a. A classical pump beam (wavelength $\simeq 775$ nm, sketched in red) injected into one or several waveguides generates photon pairs at telecom wavelength ($\simeq 1550$ nm, shown in blue) by SPDC, thanks to the strong $\chi^{(2)}$ nonlinearity of the material. These photon pairs can continuously tunnel from one waveguide to the other during their propagation, implementing random quantum walks. Compared to quantum walks in passive circuits [18, 19, 20], the walkers are here generated directly inside the device and the generation can take place at any position along the propagation axis [27]. Besides a gain of integration, this configuration allows for a progressive buildup of spatial entanglement, due to the interference between quantum walks initiated at all possible longitudinal positions, as we shall detail in the following. The device is designed so that the transverse coupling of the pump beam is negligible, so that the pump remains confined in the initially pumped waveguide(s) during propagation [35].
The precise articulation between quantum states produced by linear waveguide arrays (i.e. injected with photons produced externally) and nonlinear arrays (where photon pairs are generated continuously by internal SPDC) can be determined exactly through analytical calculations. Let |$\ket{\Psi}\!=\!\sum_{n_{s},n_{i}}\Psi(n_{s},n_{i})\ket{n_{s},n_{i}}$ denote the biphoton state generated by a nonlinear array, where $\Psi(n_{s},n_{i})$ gives the probability amplitude to detect one photon in waveguide number $n_{s}$ and the other photon in waveguide $n_{i}$. If the nonlinear array is pumped by a monochromatic pump field tuned to the single-waveguide phase matching condition, generating photon pairs by frequency-degenerate SPDC, one can show [35] that
[…]
where $\gamma$ accounts for the efficiency of the SPDC process, $z$ is the longitudinal position, $L$ is the array length, $A_{n}$ is the pump field amplitude in waveguide $n$, and $\phi_{n}(n_{s},n_{i},z)$ is the quantum state that would be produced by a linear array with two photons injected in waveguide $n$, after a propagation distance $z$ [35]. The latter state reads
$\phi_{n}(n_{s},n_{i},z)=i^{n_{s}+n_{i}-2n}\ J_{n_{s}-n}(2Cz)\ J_{n_{i}-n}(2Cz)$ (2)
where $J_{n}$ is the first-order Bessel function of index $n$ and $C$ is the coupling constant between waveguides [16].
[…]
To illustrate this interplay between quantum walks and nonlinearity, let us compare (i) the output state of a linear array injected with a path-entangled state (|$(\ket{00}+\ket{11})/\sqrt{2}$ (corresponding to two indistinguishable photons injected in either the central waveguide $n\!=\!0$ or the neighboring waveguide $n\!\!=\!\!1$), and (ii) the state generated by a nonlinear array when pumping these two waveguides in-phase ($A_{0}\!=\!A_{1}$). The corresponding intensity correlation maps are plotted in Fig. 1d and e, respectively, for $CL=5$. In the linear case (i), photons are distributed over a square-like pattern, with dominant antidiagonal lobes resulting in an antibunched state. These lobes, located near ±$\pm 2CL$, correspond to a ballistic-like propagation of both photons in opposite directions [16]. As the propagation length increases, this overall pattern expands linearly, with the biphoton wavefunction spreading over more waveguides. However the entanglement level of the state, quantified by the Schmidt number $K$ (Fig. 1f, blue), remains fixed to that of the injected state ($K\!=\!2$) since propagation in a linear optical circuit cannot increase entanglement [36].
Now turning to the nonlinear case (ii), the output state (Fig. 1e) results from the interference between linear-array states (i) obtained for continuously varying propagation lengths (Eq. (1)). This gives rise to a sharp interference pattern, where essentially only antidiagonal points survive. As the propagation length increases, the Schmidt number builds up monotonically (Fig. 1f, red). Obtaining such high-dimensional entangled states from a linear array would necessitate the delicate preparation of a quantum state of same $K$ at the input. In nonlinear arrays by contrast, the constructive interplay between quantum walks and SPDC generation provides a compact source of spatial entanglement, which can be entirely controlled from simple classical resources, namely from the spatial distribution of the input pump laser. In this work, we experimentally demonstrate this concept using a double-pump configuration.
@@
The investigated sample consists in an array of 31 coupled AlGaAs waveguides of length $L=2$ mm (Fig. 1b). The waveguides are 1.7 $\mu$m-wide and separated by 450 nm gaps, resulting in a coupling constant $C_{\rm TE}=2.7$ mm-1 in TE polarization and $C_{\rm TM}=2.4$ mm-1 in TM polarization. The 7 central waveguides of the array are connected to fan-in injection waveguides and fan-out collection waveguides to facilitate their individual optical addressing (Fig. 1c). These injection and collection waveguides have a larger width (6 $\mu$m) to shift their nonlinear resonance wavelength so that the injected pump laser generates photon pairs only in the central array region, exploiting a modal phase-matching scheme [35, 37, 38, 39, 40].
@@
The experiments are carried out by injecting a TE-polarized continuous-wave pump laser with a microscope objective into one or several waveguides (depending on the targeted quantum state), to generate orthogonally polarized signal and idler photons by type-2 SPDC. These SPDC photons undergo random quantum walks and are collected at the output of the fan-out waveguides using a commercial lensed fiber array connected to superconducting nanowire single-photon detectors (SNSPD). Correlation events, measured by a time tagger, yield the spatial correlation matrix $\Gamma_{n_{s},n_{i}}$.
@@
Figure 2: (a) Biphoton correlations measured in the central waveguide ($\Gamma_{0,0}$, black), first diagonal ($\Gamma_{1,1}+\Gamma_{-1,-1}$, red) and first antidiagonal ($\Gamma_{1,-1}+\Gamma_{-1,1}$, blue) points of the correlation matrix when pumping the $n=0$ waveguide (raw coincidence counts). (b-c-d) Measured correlation matrices when pumping (b) the central waveguide $n=0$, (c) guides $n=0$ and $n=1$ in phase and (d) guides $n=0$ and $n=1$ out of phase. (e) Cross-correlation term $\Gamma_{0,1}$ measured for varying phase offset $\varphi$ between the two pumped waveguides ($n=0$ and $1$); circles show raw coincidence counts with Poissonian error bar, crosses show noise counts, and the black line is a sinusoidal fit. (f-g-h) Numerically simulated correlation matrices corresponding to the experimental cases (b-c-d).
@@
The pump laser is first injected into the central waveguide ($n=0$) of the array, and coincidences are monitored as a function of the laser wavelength for a coupled pump power of 1 mW. Here and in the subsequent measurements, a bandpass filter is used before photon detection to reduce the emission bandwidth to 16 nm and effectively suppress any possible coupling between the spatial and spectral properties of the biphoton states [35]. As shown in Fig. 2a, the coincidences in the central waveguide ($\Gamma_{0,0}$, black points) show a resonance centered at $\lambda_{p,0}=783.45$ nm, which corresponds (up to 0.05 nm) to the resonance wavelength which we measured for a single (uncoupled) waveguide. The diagonal coincidences ($\Gamma_{1,1}+\Gamma_{-1,-1}$, red points) and antidiagonal coincidences ($\Gamma_{1,-1}+\Gamma_{-1,1}$, blue points) display a similar resonant behavior.
We next set the pump wavelength to $\lambda_{p,0}$ and proceed to the measurement of the full correlation matrix within the Hilbert space formed by the 5 central waveguides. The result is shown in Fig. 2b and compared to numerical simulations in Fig. 2f. As anticipated by the results of Fig. 2a, we observe strong and selective correlations along the diagonal and the antidiagonal axes. That is, photons have an enhanced probability to exit the device either through the same waveguide ($n_{s}=n_{i}$, spatial bunching) or through opposite waveguides ($n_{s}=-n_{i}$, spatial antibunching). The non-classicality of the measured correlations can be quantified by using the criteria derived by Bromberg et al. [16, 18]. For classical light, the diagonal and off-diagonal correlations must satisfy the inequality $\Gamma_{n_{i},n_{s}}\geq\frac{1}{3}\sqrt{\Gamma_{n_{i},n_{i}}\Gamma_{n_{s},n_{s}}}$. For the measured state of Fig. 2b, this inequality is violated (up to 10 standard deviations) by several points of the measured correlation matrix [35], indicating spatial correlations that cannot be generated by classical means. The overlap between the experimental ($\Gamma_{n_{s},n_{i}}^{\rm exp}$) and theoretical ($\Gamma_{n_{s},n_{i}}^{\rm th}$) correlations matrices can be quantified using the similarity parameter [18, 19] $S=\big{(}\sum\sqrt{\Gamma_{n_{s},n_{i}}^{\rm exp}\Gamma_{n_{s},n_{i}}^{\rm th}}\big{)}^{2}/\left(\sum\Gamma_{n_{s},n_{i}}^{\rm exp}\sum\Gamma_{n_{s},n_{i}}^{\rm th}\right)$. We obtain $S=98.4\pm 1.1\%$ (without adjustable parameter), illustrating the high fidelity of the source. We note that the simulation of Fig. 2f is exact, i.e. it takes into account the finite number of waveguides, the finite bandwidth of the photons pairs and the slight polarization and frequency-dependence of the coupling constant; it is however very close to the result of Eq. (1) (mutual similarity $98.8\%$), validating the relevance of our analytical model. The measured correlation map is robust to the pump power; as the chip length increases, correlations spread across farther waveguides but the overall pattern remains qualitatively unchanged (after a short stabilization regime) [35].
@@
We then set $\varphi=\pi$, corresponding to the $n=0$ and $n=1$ waveguides pumped in phase opposition. The measured correlation matrix is shown in Fig. 2d along with simulation results in 2h (similarity $96.9\pm 1.0\%$). In this case, we observe that spatial bunching is selectively enhanced, while the spreading of correlations is comparable to the single-pump configuration. However, the total SPDC signal (integrated over all measured points of the correlation matrix) is stronger (by a factor $\simeq$ 2.8) for $\varphi=\pi$ than for $\varphi=0$, despite the fact that the input pump power is the same. As the transverse coupling of the pump beam is negligible, the two pump beams do not interfere classically and this interference effect is purely quantum in nature: the biphoton states generated in the two pumped waveguides interfere constructively when $\varphi=\pi$ and destructively when $\varphi=0$, and it can be shown analytically [35] that the total SPDC intensity is modulated as
@@
This intriguing behavior, with a constructive interference occurring for out-of-phase pumping, can be understood intuitively in the following manner. In the single-pump configuration, correlations are mainly located on the diagonal and antidiagonal (Figs. 2b,f). According to coupled-mode theory, a single photon acquires a phase of $\pi/2$ when it tunnels to a neighboring waveguide; hence, starting from its generation in the central waveguide $n=0$, the biphoton state acquires a phase $\pi$ when it moves one step along the diagonal or the antidiagonal, as this requires two single-photon jumps (either on the same or on opposite sides). The phase of the wavefunction thus changes by $\pi$ between each two successive points along the diagonal and antidiagonal. Now, as shown by Eq. (1), when pumping two neighboring waveguides $n=0$ and $1$, an interference occurs between the biphoton states |$\ket{\Psi}_{n=0}$ and |$\ket{\Psi}_{n=1}$ generated in these two waveguides. When the two pumps are in phase (resp. out of phase), this interference is destructive (resp. constructive) on the diagonal points (which are aligned for the two states) because of the $\pi$ phase alternation structure described above, in good agreement with the experiments and simulations. This interference is only partial since the two biphoton states have different amplitudes on the interfering points; however it plays a determinant role in the total SPDC efficiency since an important part of the signal is concentrated in each pumped waveguide, providing an intuitive argument for the total SPDC intensity being maximum when $\varphi=\pi$, in agreement with Eq. (4). The situation is reversed in the case of antidiagonal points. However, on the antidiagonal passing through $(0,1)$ the two biphoton states interfere with the exact same amplitude, leading to a perfect interference and thus a high spatial spreading of the correlations when $\varphi=0$, in good agreement with Figs. 2c and 2g. This demonstrated quantum interference effect provides the building block for engineering spatially entangled states in continuously-coupled nonlinear waveguides.
@@
In summary, we have demonstrated a compact and versatile source of spatial entanglement based on quantum walks in semiconductor AlGaAs nonlinear waveguide arrays. Tailoring the spatial profile of the pump field allows to reconfigure the output quantum state and implement various types of spatial correlations, well accounted for by analytical calculations and numerical simulations. Building upon these results, a wider zoology of quantum states could be produced by pumping more waveguides, with controlled intensity and phase relation between them, via on-chip phase-shifters exploiting the strong electro-optic effect of AlGaAs [41, 40]. Pushing further the integration, the possibility to integrate the pump laser directly within the nonlinear medium [39] constitutes a distinctive asset of the AlGaAs platform compared to dielectric and silicon-based photonic circuits where spatially entangled states have been studied previously [9].
@@
In addition, the strong second-order nonlinearity of AlGaAs could be leveraged to push nonlinear waveguide arrays into the squeezing regime [42, 43], enabling the generation of highly multimode squeezed states in a flexible manner [32, 33, 34]. Finally, other lattice geometries can be implemented by modifying the propagation and coupling constants of the waveguides either at the fabrication step (by tuning the widths and distances between waveguides) or in-situ by exploiting the electro-optic effect of AlGaAs. This would enrich the possibilities of quantum state engineering to implement e.g. biphoton W-states [44] or discrete fractional Fourier transforms [45] of entangled states. The realization of quasi-periodic geometries (e.g. Fibonacci and Aubry-André) would allow investigating topological effects in the quantum regime [24] or the Anderson localization of multi-particle states with peculiar cascaded or reentrant behaviors [46, 47], making the demonstrated platform appealing to simulate in a controlled environment physical problems otherwise difficult to access in condensed-matter systems.